32. Canadian law prevents the sale of fish for consumption containing more that 2.00 ppm of PCBs. A 227-g sample of fish is tested and found to contain the maximum allowable concentration of PCBs. Determine the mass of PCBs in this sample of fish.


33. It is considered unsafe to have more than 50.0 ppm of arsenic in drinking water. If you have a bottle of water containing 250 g of water with this level of arsenic, what mass of arsenic would you ingest?


1
Expert's answer
2021-03-16T08:38:05-0400

32. 1000 ppm means that there is 1g of a substance in 1000 g of a sample. If the concentration is 2.00 ppm, there are 0.002 g of PCBs in 1000 g. So, in 227 g:

m=227*0.002/1000=0.00045(g).


33. If the concentration is 50.0 ppm, there are 0.05 g of arsenic in 1000 g of water. So, in 250 g:

m=250*0.05/1000=0.0125(g).